OpenStudy (anonymous):

What is the slope of the line that contains the points (12, 17) and (-8, -17)? Select one of the options below as your answer: A. 0 B. 10/17 C. 17/10 D. 170/2

OpenStudy (pounce0129):

use the formula m= y2-y1/x2-x1

OpenStudy (anonymous):

use point-slope formula \[(y_2-y_1)=m(x_2-x_1)\] where (12,17) and (-8, -17) are (x1,y1) and (x2,y2) that will tell you m (the slope) then use \[(y-y_1)=m(x-x_1)\] and plug in your slope and one of the points then solve for y